OpenStudy (anonymous):

Find the value of k so that the ordered pair satisfies the equation: 6x-ky=k (2,2) k=6? k=9? k=4? k=-4?

OpenStudy (anonymous):

k = 4

OpenStudy (anonymous):

6x - ky = k 6x = ky + k 6x = k (y + 1) k = 6x/(y + 1) and from the givent point (2, 2), we know that x = 2 & y =2, after putting values of x & y, we get k = 6*2/(2 + 1) k = 12/3 k = 4 Answer :) Hope it helps

OpenStudy (anonymous):

thank you so much!!!